What are the length and width of a rectangle if the area is 13.5 square feet and the length of one side is .2 the measure of the perimeter?
Answer by venugopalramana(3286)

let length of rectangle be L and width be W.
area = L*W=13.5..........(1)
perimeter = 2*(L+W)
length of one side =L = 0.2 times perimeter=0.2*2*(L+W)
L=0.4*L+0.4*W
0.6*L=0.4*W
1.5*L = W....substituting in (1)we get
1.5*L*L=13.5
L*L= 13.5/1.5 =9
L=3
hence W = 13.5/3 = 4.5